OpenStudy (drewbooks):

Write (1/3i) - (-6 + 2/3i) as a complex number in standard form.

OpenStudy (drewbooks):

To clarify: \[(\frac{ 1 }{ 3 }i) - (-6 + \frac{ 2 }{ 3 }i)\]

OpenStudy (solomonzelman):

you need the a+bi form

OpenStudy (welshfella):

just simplify the imaginary parts (containing the 'i')

OpenStudy (solomonzelman):

-(-C+D) --> C-D

OpenStudy (solomonzelman):

expand the parenthesis and add/subtract

OpenStudy (drewbooks):

So the -(-6+2/3i) is 6-2/3i?

OpenStudy (solomonzelman):

yes, and than add (1/3)i to it

OpenStudy (drewbooks):

(6-1/3i)?

OpenStudy (solomonzelman):

yes, close, actually, you need a+bi form, so \(\large\color{black}{ \displaystyle 6+\left(-\frac{1}{3}\right){\rm i}}\)